# Heuristic search for linear positive systems

### v1

@@ 5 Heuristic search for positive systems
$x(t+1)=(A_{S}+BK)x(t),\;\;\;K\in\overline{\mathcal{K}}_{S}\;\textnormal{or}\;\underline{\mathcal{K}}_{S}$
This construction preserves any quantity leaving the search area $S$ by transfer to some state $x^{i}$, $i\in\mathcal{V}\setminus S$ in that state, no longer affecting the states in $S$.
Let $\mathcal{I}_{S}$ denote a matrix with zeros in all elements except diagonal entries 1 with indices corresponding to states in $S$. Define the upper bound local optimum $\overline{g}_{S}$ as the solution to
$\overline{g}_{S}=\mathcal{I}_{S}s+A_{S}^{\top}\overline{g}_{S}+\mathcal{I}_{S}\min_{K\in\overline{\mathcal{K}}_{S}}\left[K^{\top}(r+B^{\top}\overline{g}_{S})\right]$ (12)
with ${\overline{g}^{i}_{S}=\overline{h}^{i}}$ for ${i\in\mathcal{V}\setminus S}$. The lower bound local optimum $\underline{g}_{S}$ is defined analogously using ${\underline{g}_{S}^{i}=\underline{h}^{i}}$ and ${K\in\mathcal{\underline{K}}_{S}}$. With this definition we are ready to state the following heuristic search algorithm.
Algorithm 1 Heuristic search
Inputs: $x_{0}$, $\hat{K}$, $A$, $B$, $E$, $s$, $r$, $\gamma$
1:define $S\leftarrow\{i:x^{i}_{0}>0\}$
@@
5:     extract $\overline{K}_{S}=\textnormal{arg}\min_{K\in\mathcal{K}_{S}}K^{\top}(r+B^{\top}\overline{g}_{S})$ and $\underline{K}_{S}=\textnormal{arg}\min_{K\in\mathcal{K}_{S}}K^{\top}(r+B^{\top}\underline{g}_{S})$
6:     calculate ${\overline{x}=\lim_{k\rightarrow\infty}(A_{S}+B\overline{K}_{S})^{k}x_{0}}$ and ${\underline{x}=\lim_{k\rightarrow\infty}(A_{S}+B\underline{K}_{S})^{k}x_{0}}$
7:     set $S\leftarrow S\cup\textnormal{arg}\max\limits_{i\in\mathcal{V}\setminus S}\left[(\overline{h}^{i}-\underline{h}^{i})(\overline{x}^{i}+\underline{x}^{i})\right]$
@@
The algorithm terminates when the bounds provided by the heuristic function can guarantee that the worst-case cost $\overline{g}^{\top}x_{0}$ is at most $\gamma$ times larger than the optimal $p^{\top}x_{0}$.
@@
Algorithm 1 with $\gamma=1$ will terminate in finite time and return the optimal cost function in all states for which $x_{0}$ is non-zero.
@@
We first prove that the algorithm terminates. Note that step 4 can be solved by means of linear programming as a direct result of Theorem 1, and can thus be completed in finite time. With a view towards generating a contradiction, assume that the algorithm does not terminate. By iteration of step 7, we must then eventually have $S=\mathcal{V}$. This means that step 4 solves (12) for the full state, which is equivalent to (2), yielding ${\overline{g}_{S}=\underline{g}_{S}=p}$ according to Theorem 1. This in turn means that ${\overline{g}_{S}^{\top}x_{0}/\underline{g}_{S}^{\top}x_{0}=1}$, and the algorithm will terminate in the next step, contradicting our assumption.
@@
Initialization of $\overline{g}_{S}$ and $\underline{g}_{S}$ in accordance with Proposition 4.3 means that ${\underline{g}_{S}\leq p\leq\overline{g}_{S}}$ holds initially. This relation is also preserved under iteration of Algorithm 1, as the operation in step 4 solves (12), which has a strictly larger (in the case of $\overline{g}_{S}$) solution than the optimal cost of the full system, meaning that ${p\leq\overline{g}_{S}}$ for all iterates. A symmetrical result holds for $\underline{g}_{S}$. Since ${x_{0}\geq 0}$, the only possible solutions fulfilling (13) are those for which ${\underline{g}_{S}^{\top}x_{0}=p^{\top}x_{0}=\overline{g}_{S}^{\top}x_{0}}$. Consequently, we can only have ${\overline{g}_{S}^{i}\neq p^{i}}$ and ${\underline{g}_{S}^{i}\neq p^{i}}$ for elements $i$ such that ${x_{0}^{i}=0}$. $\blacksquare$
@@
where the cost is attained by following the worst-case policy associated with $\overline{g}_{S}$. $\blacksquare$ Step 4 of Algorithm 1 can, for example, be performed by fixed point iteration of (12), or by solving a linear program like ($iii$) in Theorem 1. Either way, the complexity of the calculation scales with the size of the local search area $S$, as opposed to the full state space. In step 5, only local information of $r$ and $B$ is required, as all other elements are multiplied by zero due to the definition of the set $\mathcal{K}_{S}$. Additionally, if step 4 is performed by e.g. Value Iteration or Policy Iteration, the policies $\overline{K}_{S}$ and $\underline{K}_{S}$ are naturally obtained, removing the need to perform step 5.
The monotone convergence of Value Iteration means that we can leverage the previous iterations of $\overline{g}_{S}$ and $\underline{g}_{S}$ as starting points when performing step 4. In light of Proposition 4.5, this guarantees a faster rate of convergence for Value Iteration compared to starting from the initial heuristics of Proposition 4.3. On the other hand, if step 4 is performed by solving a linear program on the form in ($iii$) of Theorem 1, the monotone improvement property means that each iterate of $\overline{g}_{S}$ and $\underline{g}_{S}$ is closer to the solution than the previous. Meanwhile, as noted previously, the conditions for a consistent heuristic are identical to the constraints of the linear program. For the upper bound heuristic the objective in ($iii$) can be minimized with the constraint inequality for $p$ reversed, resulting in a linear program with identical solution where an intermediate upper bound $\overline{g}_{S}$ satisfies the constraints.
We also note that the optimization over local controllers in step 5 can be simplified in the event that the expression
@@
for some index $i$ yields the same result for all $g$ in the interval $\underline{g}_{S}\leq g\leq\overline{g}_{S}$. In this case, one action is optimal for all possible values of $p$ within the current bounds. This action can thus be fixed in all future iterations, reducing the computational complexity. This idea of "locking in" actions that can be certified as optimal within the given bounds is leveraged in several heuristic search frameworks for SSP, see e.g. the Hierarchical Dirichlet Process (HDP) algorithm in (Bonet & Geffner, 2003).
@@
Step 6 is significantly reduced in size for systems with sparse dynamics. The states $\overline{x}$ and $\underline{x}$ are only calculated for states that are directly impacted by those in the search area $S$. The priority of states outside $S$ in step 7 is not uniquely optimal, see (McMahan et al., 2005) for a discussion of different methods. The current setting differs in that dynamics in continuous state are simulated to determine which states have the most potential to improve the total cost, both as a consequence of a large uncertainty $\overline{h^{i}}-\underline{h}^{i}$ and as a result of a high degree of utilization $\overline{x}^{i}+\underline{x}^{i}$ in either the worst or best case. Changing or weighting this decision offers the chance to tailor the algorithm for, e.g., more optimistic or pessimistic intermediate solutions.
Figure 4: Upper and lower bound of the total cost in Example 6.1. The red lines highlight the iterations for which the state is displayed in more detail in Figure 5. The value $\gamma$ specifies the stopping condition in Algorithm 1 and guarantees a performance of at least $\overline{g}^{\top}x_{0}\leq\gamma p^{\top}x_{0}$ when the algorithm terminates.

### v2

@@ 5.1 Stabilized systems
$x(t+1)=(A_{S}+BK)x(t),\;\;\;\overline{\mathcal{K}}_{S}\;\textnormal{or}\;\underline{\mathcal{K}}_{S}.$
This construction preserves any quantity leaving the search area $S$ by transfer to some state $x_{i}$, $i\in\mathcal{V}\setminus S$ in that state, no longer affecting the states in $S$. The final distribution is then used in the prioritization of states when expanding the search area.
Let $I_{S}$ denote a matrix with zeros in all elements except diagonal entries 1 at indices corresponding to states in $S$. Define the upper bound local optimum $\overline{g}_{S}\in\mathbb{R}^{|S|}_{+}$ as the solution to
$\overline{g}_{S}=I_{S}s+A_{S}^{\top}\overline{g}_{S}+I_{S}\min_{K\in{\overline{\mathcal{K}}_{S}}}\left[K^{\top}(r+B^{\top}\overline{g}_{S})\right]$ (15)
with $(\overline{g}_{S})_{i}=\overline{h}_{i}$} for $i\in\mathcal{V}\setminus S$}. The lower bound local optimum $\underline{g}_{S}$ is defined analogously using $(\underline{g}_{S})_{i}=\underline{h}_{i}$} and $K\in\overline{\mathcal{K}}_{S}$}. With these definition we are ready to state Algorithm 1. The algorithm terminates when the bounds provided by the heuristic function can guarantee that the worst-case cost $\overline{g}^{\top}x_{0}$ is at most $\gamma$ times larger than the optimal $p^{\top}x_{0}$.
Algorithm 1 Heuristic search with initial policy
Inputs: $x_{0}$, $A$, $B$, $E$, $s$, $r$, $\gamma$, $\hat{K}$
1:define $S\leftarrow\{i:(x_{0})_{i}>0\}$
@@
5:   extract $\overline{K}_{S}=\textnormal{arg}\min_{K\in\overline{\mathcal{K}}_{S}}K^{\top}(r+B^{\top}\overline{g}_{S})$
6:   calculate $\overline{x}=\lim_{k\rightarrow\infty}(A_{S}+B\overline{K}_{S})^{k}x_{0}$}
7:  set $S\leftarrow S\cup\textnormal{arg}\max\limits_{i\in\mathcal{V}\setminus S}\left[(\overline{h}_{i}-\underline{h}_{i}){\overline{x}_{i}}\right]$
@@
Consider a problem on the form (1) with finite value for $x_{0}\in\mathbb{R}^{n}_{+}$. Algorithm 1 with $\gamma=1$ will terminate in finite time and return the optimal cost function in all states for which $x_{0}$ is non-zero.
@@
We first prove that the algorithm terminates. Note that step 4 can be solved by means of linear programming as a direct result of Theorem 1, and can thus be completed in finite time. With a view towards generating a contradiction, assume that the algorithm does not terminate. By iteration of step 7, we must then eventually have $S=\mathcal{V}$. This means that step 4 solves (15) for the full state, which is equivalent to (3), yielding $\overline{g}_{S}=\underline{g}_{S}=p$} according to Theorem 1, since the full problem has a finite optimal value. This in turn means that $\overline{g}_{S}^{\top}x_{0}/\underline{g}_{S}^{\top}x_{0}=1$}, and the algorithm will terminate in the next step, contradicting our assumption.
@@
Initialization of $\overline{g}_{S}$ and $\underline{g}_{S}$ in accordance with Proposition 4.2 means that $\underline{g}_{S}\leq p\leq\overline{g}_{S}$} holds initially. This relation is also preserved under iteration of Algorithm 1, as the operation in step 4 solves (15), which has a strictly larger (in the case of $\overline{g}_{S}$) solution than the optimal cost of the full system, meaning that $p\leq\overline{g}_{S}$} for all iterates. A symmetrical result holds for $\underline{g}_{S}$. Since $x_{0}\geq 0$}, the only possible solutions fulfilling (16) are those for which $\underline{g}_{S}^{\top}x_{0}=p^{\top}x_{0}=\overline{g}_{S}^{\top}x_{0}$}. Consequently, we can only have $(\overline{g}_{S})_{i}\neq p_{i}$} and $(\underline{g}_{S})_{i}\neq p_{i}$} for elements $i$ such that $(x_{0})_{i}=0$}. $\blacksquare$
@@
where the cost is attained by following the worst-case policy associated with $\overline{g}_{S}$. $\blacksquare$
Step 4 of Algorithm 1 can, for example, be performed by fixed point iteration of (15), or by solving a linear program like (iiiiii) in Theorem 1. Either way, the complexity of the calculation scales with the size of the local search area $S$, as opposed to the full state space. In step 5, only local information of $r$ and $B$ is required, as all other elements are multiplied by zero due to the definition of the set $\mathcal{K}_{S}$. Additionally, if step 4 is performed by e.g. value iteration or policy iteration, the policy $\overline{K}_{S}$ is naturally obtained, removing the need to perform step 5.
As a consequence of Proposition 4.3, we can show that each iteration and expansion of the search space $S$ monotonically improves the cost. Meanwhile, as noted previously, the conditions for a consistent heuristic are identical to the constraints of the linear program. This means that, when solving step 4 using linear programming, the previous solution is a feasible point from which we can initialize the program of the next iteration.
Proposition 5.3.
Given a stabilizing policy $\hat{K}$, Algorithm 1 converges monotonically.
{pf}
Let $S_{1}$ and $S_{2}$ be the sets corresponding to the search area in subsequent iterations of Algorithm 1. Consequently, $S_{1}\subset S_{2}$. When evaluating (15) for $S_{2}$, we see that the contribution to $\overline{g}_{S_{2}}$ from states outside $S_{2}$ is identical to the case of (15) for $S_{1}$. The only difference, then, is the influence of the state, $i$ say, added to the search area between the two iterations. This influence is chosen as $\min\{r_{i}+B_{i}^{\top}\overline{g}_{S_{2}}\}E_{i}$, rather than being fixed to the previous suboptimal policy. We now have that $\overline{g}_{S_{1}}\leq\overline{h}$} implies $\overline{g}_{S_{2}}\leq\overline{g}_{S_{1}}$}. The first step of the induction follows immediately from consistency of $\overline{h}$, as given in Proposition 4.2. $\blacksquare$
We note that the optimization over local controllers in step 5 can be simplified in the event that the expression
@@
for some index $i$ yields the same result for all $g$ in the interval $\underline{g}_{S}\leq g\leq\overline{g}_{S}$. In this case, one action is optimal for all possible values of $p$ within the current bounds. This action can thus be fixed in all future iterations, reducing the computational complexity. This idea of "locking in" actions that can be certified as optimal within the given bounds is leveraged in several heuristic search frameworks for SSP, see e.g. the Hierarchical Dirichlet Process (HDP) algorithm in (Bonet and Geffner, 2003).
@@
Step 6 is significantly reduced in size for systems with sparse dynamics. The state $\overline{x}$ is only calculated for states that are directly impacted by those in the search area $S$. The priority of states outside $S$ in step 7 is not uniquely optimal, see (McMahan et al., 2005) for a discussion of different methods. The current setting differs in that dynamics in continuous state are simulated to determine which states have the most potential to improve the total cost, both as a consequence of a large uncertainty $\overline{h_{i}}-\underline{h}_{i}$ and as a result of a high degree of utilization $\overline{x}_{i}$ in the worst case. Changing or weighting this decision offers the chance to tailor the algorithm for, e.g., more optimistic or pessimistic intermediate solutions.
Remark 5.4.
In the case $E=I$, the heuristics contain all information required to find the optimal policy for a subset $S$ of nodes. In other words, we can find the optimal parameter of (3) by solving the smaller subproblem (15) with local policy $K\in\overline{\mathcal{K}}_{S}$ when the optimal solution is known for states outside $S$. This is no longer the case when $E\neq I$, as the sum in the expression (3) then indicates that the optimal policy in states outside $S$ influence the optimal value. As a consequence, we expect deteriorated performance of the heuristic search for $E$ with large off-diagonal elements. This serves to illustrate the difference between the class of SSP and the full class (1), and explains the remarkable success of heuristic methods for SSP in particular.
5.2 Unstable systems
In situations where no initial stabilizing policy is known, heuristic bounds can still help guide the search for an optimal policy. However, the method presented above must be modified. Consider the following implicit equation for the upper bound solution on the search area $S$:
$\overline{g}_{S}=I_{S}s+A_{S}^{\top}\overline{g}_{S}+I_{S}\min_{K\in{\underline{\mathcal{K}}_{S}}}\left[K^{\top}(r+B^{\top}\overline{g}_{S})\right]$} (17)
with $\overline{g}_{i}=\overline{h}_{i}$ for $i\in\mathcal{V}\setminus S$. The right-hand expression is identical to that in (15) apart from the use of the zero policy $K_{i}=\mathbf{0}$ outside the search area. This represents the lack of information about the behavior of the system outside $S$, removing the influence of the unknown inputs $u_{i}$ for $i\in\mathcal{V}\setminus S$ on the nodes in $S$. As in the previous section, an equivalent expression for $\underline{g}_{S}$ is obtained by instead using the lower bound heuristic $(\underline{g}_{S})_{i}=\underline{h}_{i}$ for states $i$ outside $S$. The procedure for heuristic search without prior knowledge of a stabilizing policy is presented in Algorithm 2. Just as above, (17) coincides with (3) for $S=\mathcal{V}$, guaranteeing convergence of Algorithm 2 within the prescribed performance bound $\gamma$. This follows directly from the same reasoning as used in the proof of Proposition 5.1 and Corollary 1.
Algorithm 2 Heuristic search without initial policy
Inputs: $x_{0}$, $A$, $B$, $E$, $s$, $r$, $\overline{h}$, $\underline{h}$, $\gamma$
1:define $S\leftarrow\{i:(x_{0})_{i}>0\}$
2:set $\underline{g}\leftarrow\underline{h}$}, $\overline{g}\leftarrow\overline{h}$}
3:while $\overline{g}_{S}^{\top}x_{0}/\underline{g}_{S}^{\top}x_{0}>\gamma$ do
4:  find $\overline{g}_{S}$ and $\underline{g}_{S}$ that solve (17)
5:   extract $\overline{K}_{S}=\textnormal{arg}\min_{K\in\underline{\mathcal{K}}_{S}}K^{\top}(r+B^{\top}\overline{g}_{S})$
6:   calculate $\overline{x}=\lim_{k\rightarrow\infty}(A_{S}+B\overline{K}_{S})^{k}x_{0}$}
7:  set $S\leftarrow S\cup\textnormal{arg}\max\limits_{i\in\mathcal{V}\setminus S}\left[(\overline{h}_{i}-\underline{h}_{i}){\overline{x}_{i}}\right]$
8:end while
Special care needs to be taken when evaluating the expansion of $S$ in step 7 of Algorithm 2. Since $\overline{h}$ is not guaranteed to be finite for all states, we let the product $(\overline{h}_{i}-\underline{h}_{i})\overline{x}_{i}$ take the value 0 for states where $\overline{x}_{i}=0$. This means that the algorithm prioritizes exploring unstable states that cannot be avoided in the worst case $\overline{x}$. Unstable states that can be avoided (giving $\overline{x}_{i}=0$) are deprioritized, as the heuristics give no information to guide the search in these states. Note that this implies that a stabilizing policy for the states in $S$ has been found when Algorithm 2 terminates, regardless of $\gamma$.
Proposition 5.5.
Given admissible heuristic bounds $\overline{h}$ and $\underline{h}$ such that $\underline{h}$ is consistent and $\overline{h}$ is superconsistent, Algorithm 2 converges monotonically.
{pf}
Let $S_{1}$ and $S_{2}$ be the sets corresponding to the search area in subsequent iterations of Algorithm 2, and let $i$ be the index of the state added between the two iterations, so that $S_{2}=S_{1}\cup\{i\}$. We wish to show that $\overline{g}_{S_{1}}\geq\overline{g}_{S_{2}}$. Since $\overline{h}$ is superconsistent, the element corresponding to the added state fulfills
$(\overline{g}_{S_{1}})_{i}$ $=\overline{h}_{i}$
$\geq s_{i}+A_{i}^{\top}\overline{h}+\min\{r_{i}+B_{i}^{\top}\overline{h},0\}E_{ii}$
$\geq s_{i}+A_{i}^{\top}\overline{h}+\sum_{j=1}^{n}\min\{r_{j}+B_{j}^{\top}\overline{g}_{S_{1}},0\}E_{ji}$
$=(\overline{g}_{S_{2}})_{i}.$
Here, the second inequality follows admissibility of $\overline{h}$ and the fact that all terms of the right-hand side sum are nonpositive. $\blacksquare$
Remark 5.6.
Note that, in order to attain the specified performance bound and guarantee stability, we implicitly assume that performance within the heuristic bounds $\overline{h}$ and $\underline{h}$ can be realized. This does not mean, however, that we necessarily know the policy in states outside $S$. The heuristics may be obtained from measurements of performance under some nominal operation of the system.
6 Distributed value iteration
Algorithms 1 and 2 leverage the constructed heuristic bounds to find close-to-optimal solutions for specific initializations $x_{0}$. To achieve this, centralized information about the dynamics is required for the growing search region. Below, we present a second algorithm that instead uses the local bound on performance in each state implied by the heuristics to distribute the value iteration algorithm. In contrast to standard Gauss-Seidel iteration used to distribute value iteration of (3) (as used in (Ohlin et al., 2024)), the use of upper and lower bounds allows for a distributed stopping condition based on the desired distance to optimality, with global performance guarantees.
Consider $n$ agents, each representing one element $x_{i}$ of the state vector. In the graphical interpretation of Figure 3, each agent corresponds to a node, locally determining the gain $K_{i}$ for the inputs $u_{i}$ and corresponding constraint. Each matrix $B_{i}$ as defined in the problem setup (Section 2) above governs the effect of agent $i$’s local actions. Define the local incidence matrix $W^{(i)}=B_{i}B_{i}^{\top}$} for each agent, and the following sets of neighbors:
${\mathcal{N}^{A}_{i}:=\{j:A_{ji}\neq 0,j\neq i\}}$
${\mathcal{N}^{E}_{i}=\{j:E_{ji}\neq 0,j\neq i\}}$
${\mathcal{N}_{i}^{B}=\{j:W^{(i)}_{ij}\neq 0,j\neq i\}.}$
Algorithm 3 successively improves the local bounds on the cost function for each agent via value iteration. Agent $i$ terminates when the local suboptimality gap is reduced below a chosen parameter $\gamma$. The global performance of the algorithm is shown in the following proposition. It proves that termination by one or more agents does not prevent the remainder from reaching the prescribed level of optimality. In other words, the suboptimality gap induced by agents terminating does not grow beyond the factor $\gamma$ as it propagates through the network, regardless of the number of agents. This performance is achieved by the resulting pessimistic policy $\overline{K}$, associated with the obtained upper bound.
Algorithm 3 Distributed value iteration
Inputs: $A$, $B$, $E$, $s$, $r$, $\overline{h}$, $\underline{h}$, $\gamma$
1:define $\mathcal{I}=\{1,...,n\}$
2:for all $i\in\mathcal{I}$ do
3:  set $\overline{g}_{i}\leftarrow\overline{h}_{i}$, $\underline{g}_{i}\leftarrow\underline{h}_{i}$, $\overline{q}_{i}\leftarrow 0$, $\underline{q}_{i}\leftarrow 0$
4:end for
5:while $\mathcal{I}\neq\emptyset$ do
6:  sample agent $i\in\mathcal{I}$
7:  receive $\overline{g}_{j}$, $\underline{g}_{j}$ from each neighbor $j\in\mathcal{N}_{i}^{A}\cup\mathcal{N}_{i}^{B}$
8:  receive $\overline{q}_{j}$, $\underline{q}_{j}$ from each neighbor $j\in\mathcal{N}_{i}^{E}$
9:  $\overline{q}_{i}\leftarrow\min\{r_{i}+B_{i}^{\top}\overline{g},0\}$, $\underline{q}_{i}\leftarrow\min\{r_{i}+B_{i}^{\top}\underline{g},0\}$
10:  $\overline{g}_{i}\leftarrow s_{i}+A_{i}^{\top}\overline{g}+\overline{q}_{i}E_{ii}+\sum\limits_{j\in\mathcal{N}_{i}^{E}}\overline{q}_{j}E_{ji}$
11:  $\underline{g}_{i}\leftarrow s_{i}+A_{i}^{\top}\underline{g}+\underline{q}_{i}E_{ii}+\sum\limits_{j\in\mathcal{N}_{i}^{E}}\underline{q}_{j}E_{ji}$
12:  if $\overline{g}_{i}/\underline{g}_{i}<\gamma$ then
13:   set $\mathcal{I}\leftarrow\mathcal{I}\setminus i$
14:  end if
15:end while
Proposition 6.1.
Consider a problem on the form (1) with finite value for $x_{0}\in\mathbb{R}^{n}_{+}$. Algorithm 3 converges to a solution $\overline{g}$ such that $\overline{g}\leq\gamma p$ for any given $\gamma\geq 1$.
{pf}
Algorithm 3 performs value iteration for $\overline{g}$ and $\underline{g}$ using the Gauss-Seidel method, which converges asymptotically to the fixed point of (3). When an agent terminates, its current (suboptimal) bounds are fixed. This means that the iterations of upper and lower bounds for the remaining system have two separate fixed points $\overline{g}^{*}$ and $\underline{g}^{*}$. To prove the stated performance bound, we will show that termination of one or more agents under the given condition cannot perturb the distance between the fixed points of the upper and lower bounds for the remaining system beyond a factor $\gamma$.
Consider the scenario where a subset $\mathcal{I}$ of agents are still iterating according to Algorithm 3, while the rest have terminated, partitioning the upper bound vector accordingly as $\overline{g}_{{}_{\mathcal{I}}}$ and $\overline{g}_{{}_{-\mathcal{I}}}$. Rewriting (5) gives the following expression for the fixed point:
$\begin{bmatrix}\overline{g}^{*}_{{}_{\mathcal{I}}}\\ \overline{g}^{*}_{{}_{-\mathcal{I}}}\end{bmatrix}=\underbrace{\begin{bmatrix}\overline{M}_{{}_{\mathcal{I},\mathcal{I}}}&\overline{M}_{{}_{\mathcal{I},-\mathcal{I}}}\\ \overline{M}_{{}_{-\mathcal{I},\mathcal{I}}}&\overline{M}_{{}_{-\mathcal{I},-\mathcal{I}}}\end{bmatrix}}_{=(A+B\overline{K})^{\top}}\begin{bmatrix}\overline{g}^{*}_{{}_{\mathcal{I}}}\\ \overline{g}^{*}_{{}_{-\mathcal{I}}}\end{bmatrix}+\underbrace{\begin{bmatrix}\overline{k}_{{}_{\mathcal{I}}}\\ \overline{k}_{{}_{-\mathcal{I}}}\end{bmatrix}}_{=s+\overline{K}^{\top}r}.$} (18)
Solving for $\overline{g}^{*}_{{}_{\mathcal{I}}}$ we get
$\overline{g}^{*}_{{}_{\mathcal{I}}}=(I-\overline{M}_{{}_{\mathcal{I},\mathcal{I}}})^{-1}\left(\overline{M}_{{}_{\mathcal{I},-\mathcal{I}}}\overline{g}^{*}_{{}_{-\mathcal{I}}}+\overline{k}_{{}_{\mathcal{I}}}\right).$} (19)
An equivalent expression holds for $\underline{g}^{*}_{{}_{\mathcal{I}}}$, with dynamics $\underline{M}=(A+B\underline{K})$. In order to prove that the difference between the perturbed fixed points is still within the specified bound, we need to show nonpositivity of the quantity
$\overline{g}^{*}_{{}_{\mathcal{I}}}-\gamma\underline{g}^{*}_{{}_{\mathcal{I}}}=\overline{g}^{*}_{{}_{\mathcal{I}}}-\gamma(I-\underline{M}_{{}_{\mathcal{I},\mathcal{I}}})^{-1}\left(\underline{M}_{{}_{\mathcal{I},-\mathcal{I}}}\underline{g}^{*}_{{}_{-\mathcal{I}}}+\underline{k}_{{}_{\mathcal{I}}}\right).$} (20)
For agents not in $\mathcal{I}$, the stopping condition implies that $\overline{g}^{*}_{{}_{-\mathcal{I}}}\leq\gamma\underline{g}^{*}_{{}_{-\mathcal{I}}}$}. Further, the matrix $\underline{M}_{{}_{\mathcal{I},\mathcal{I}}}$ is a principal submatrix of $(A+B\underline{K})^{\top}$ and thus Schur stable. This together with $\underline{M}_{{}_{\mathcal{I},\mathcal{I}}}\geq 0$ means that $I-\underline{M}_{{}_{\mathcal{I},\mathcal{I}}}$ is an M-matrix, which in turn implies that $(I-\underline{M}_{{}_{\mathcal{I},\mathcal{I}}})^{-1}\geq 0$. We therefore obtain the following element-wise upper bound on (20):
${\overline{g}^{*}_{{}_{\mathcal{I}}}-\gamma\underline{g}^{*}_{{}_{\mathcal{I}}}\leq\overline{g}^{*}_{{}_{\mathcal{I}}}-(I-\underline{M}_{{}_{\mathcal{I},\mathcal{I}}})^{-1}\left(\underline{M}_{{}_{\mathcal{I},-\mathcal{I}}}\overline{g}^{*}_{{}_{-\mathcal{I}}}+\underline{k}_{{}_{\mathcal{I}}}\right)}\\ {-(\gamma-1)(I-\underline{M}_{{}_{\mathcal{I},\mathcal{I}}})^{-1}\underline{k}_{{}_{\mathcal{I}}}.}$ (21)
Here, the terms have been rearranged to illustrate the fact that the second term corresponds to the expression (19) for the upper bound if the optimal policy is replaced by $\underline{K}$. This policy cannot outperform  $\overline{K}$, and Proposition 4.3 shows that the optimal upper bound is element-wise less than or equal to those under other policies. Thus, the sum of the first two terms in (21) is nonpositive in every element. Nonpositivity of the right-hand side in (21) now follows from $\underline{k}_{{}_{\mathcal{I}}}>0$, completing the proof. $\blacksquare$